If an obtuse, acute, and right triangle each have an area of 12 square units. What is their base and height
Found 2 solutions by richard1234, Alan3354:
Answer by richard1234(7193)

Not enough information...infinitely many triangles can have an area of 12 square units.

Answer by Alan3354(69443)

If an obtuse, acute, and right triangle each have an area of 12 square units. What is their base and height
--------------
b*h/2 = 12
b*h = 24
b = 24/h or h = 24/b
That's all that can be determined.